The width of a rectangle is 1 ft less than the length. The area is 2 ft^2. Find the length and the width.
^means it is raising to that power

Answer by algebrahouse.com(1659)

x = length
x - 1 = width {width is one less than the length}

Area of a rectangle is length x width.

x(x - 1) = 2 {area is length times width}
x² - x = 2 {used distributive property}
x² - x - 2 = 0 {subtracted 2 from both sides}
(x - 2)(x + 1) = 0 {factored into two binomials}
x - 2 = 0 or x + 1 = 0 {set each factor equal to 0}
x = 2 or x = -1 {solved each equation for x}
x = 2 {length cannot be negative}
x - 1 = 1 {substituted 2, in for x, into x - 1}

width = 1 ft and length = 2 ft

For more help from me, visit: www.algebrahouse.com